# Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding

### v2

@@ 3.1. Data
MS MARCO v1 was created from the MS MARCO reading comprehension dataset (Bajaj et al., 2016) and it has two collections: passages and documents, which are related. MS MARCO v1 comes with large query sets, which is particularly useful for training and testing models in the big-data regime. These query sets consist of question-like queries sampled from the Bing search engine log (with subsequent filtering) (Craswell et al., 2020). Note that queries are not necessarily proper English questions, e.g., “lyme disease symptoms mood”, but they are answerable by a short passage retrieved from a set of about 3.6M Web documents (Bajaj et al., 2016).
MS MARCO v1 test sets were created in two stages, where initially relevance judgements were created for the passage variant of the dataset. Then, document-level relevance labels were created by transferring passage-level relevance to original documents from which passages were extracted. To assess positional bias, we mapped relevant passages (from the MS MARCO Passage collection) to their positions in documents. Because document and passage texts were collected at different times this lead to some content divergence (Craswell et al., 2020) and made exact mapping impossible: In particular, Hofstätter et al. (Hofstätter et al., 2020) were able to match only 32% of the passages (which we think is not a sufficient coverage).
Table 4. Distribution of Start/End Positions of Relevant Passages Inside Documents
input chunk #
MS MARCO dev
FIRA (Hofstätter et al., 2020)
(crowd-sourced)
start end start end
1 85.9% 71.0% 83.8% 76.4%
2 9.1% 14.9% 9.9% 15.3%
3 2.6% 6.1% 2.3% 3.9%
4 1.2% 3.0% 2.2% 2.2%
5 0.6% 1.4% 0.7% 0.9%
6 0.6% 1.2% 0.4% 0.5%
6+ 0.1% 2.5% 0.7% 0.7%
  • Chunk size is 477 BERT tokens.
To obtain a more comprehensive mapping we resorted to approximate matching and were able to match about 85% of the passages. We manually inspected a sample of matched passages to ensure that the matching procedure was reliable. Moreover, the distribution of positions of relevant passages matches that of a related FIRA dataset (Hofstätter et al., 2020), where such information was collected by crowdsourcing. Positional bias information is summarized in Table 4.
Relevance labels in the training and development sets are “sparse”: There is about one positive example per query without explicit negatives. In addition to sparse relevance judgements—separated into training and developments subsets—there is a small number (98) of queries that have “dense” judgements provided by NIST assessors for TREC 2019 and 2020 deep learning (DL) tracks (Craswell et al., 2020).
@@
The MS MARCO FarRelevant collection was created from the MS MARCO passage collection in such a way that each document contained exactly one relevant passage and this passage did not start before token 512. We created one document for each training or development query that had a relevant passage. Note that we generated about 7K test and close to 500K training queries, but used only 50K for fine-tuning and 1K for testing. On one hand, this was sufficient for accurate training and testing and, on the other hand, it reduced experimentation time and cost.
More specifically the dataset was created as follows: Assume that $C_{t}$ is the number of tokens in the passage:
  • Select randomly a document length between $512+C_{t}$ and 1431;
  • Using rejection sampling, obtain $K_{1}$ non-relevant samples such that their length exceeds 512, but the length of $K_{1}-1$ first samples is at most 512.
  • Discard the last sampled passage and randomly mix the last $K_{2}$ non-relevant passages with a single relevant passage.
Table 5. Comparison of Long-context Models to Respective FirstP baselines and Prior Art.
Model MS MARCO TREC DL Robust04
dev 2019 2020 2021 title description
@@
Prior work (FirstP, MaxP), Zhang et al. (Zhang et al., 2021)
FirstP (BERT) (Zhang et al., 2021) 0.449 0.510
MaxP (BERT) (Zhang et al., 2021) 0.477 $(+6.2\%)$ 0.530 $(+3.9\%)$
MaxP (ELECTRA) (Zhang et al., 2021) 0.523 0.574
Prior work (PARADE) Li et al. (Li et al., 2024)
PARADE Attn (ELECTRA) (Li et al., 2024) 0.527 0.587
PARADE Max (ELECTRA) (Li et al., 2024) 0.679 0.613 0.544 0.602
PARADE Transf-RAND (ELECTRA) (Li et al., 2024) 0.650 0.601 0.566 0.613
Our results
FirstP (BERT) 0.394 0.631 0.598 0.660 0.475 0.527
FirstP (ELECTRA) 0.417 0.652 0.642 0.686 0.492 0.552
FirstP (DEBERTA) 0.415 0.675 0.629 0.702 0.534 0.596
FirstP (Longformer) 0.404 0.657 0.616 0.654 0.483 0.540
FirstP (Big-Bird) 0.408 0.663 0.620 0.679 0.507 0.560
FirstP (JINA) 0.422 0.658 0.618 0.679 0.488 0.532
FirstP (MOSAIC) 0.423 0.654 0.607 0.662 0.453 0.538
3.2. Setup
In our setup, a ranker is applied to the output of the first-stage retrieval model, also known as a candidate-generator. Thus, the ranker is trained to distinguish between positive examples (known relevant documents) and hard negative examples (documents not marked as relevant) sampled from the set of top-$k$ candidates returned by the candidate generator. We used $k=100$ for MS MARCO and MS MARCO FarRelevant and $k=1000$ for Robust04.
Except for ablations, each model was trained using three seeds. To compute statistical significance, we averaged query-specific metric values over these seeds. To enable efficient training and evaluation of the large number of models, documents were truncated to have at most 1431 BERT tokens. Queries were padded to 32 BERT tokens with padding being masked out during training (longer queries were truncated). Thus, for SplitP models with greedy partitioning into disjoint chunks, long document were split into at most three chunks containing 477 document tokens (each concatenated with up to 32 query tokens plus three special tokens). Except for the aggregation and interaction Transformers, we use a base, i.e., a 12-layer Transformer (Vaswani et al., 2017) models. Further model- and training-related details are available in the online appendix (onl, [n. d.]).
Depending on the collection we computed a subset of the following metrics: the mean reciprocal rank (MRR), the non-discounted cumulative gain at rank $k$ (NDCG@K) (Järvelin and Kekäläinen, 2002), the mean average precision (MAP), and precision at rank (P@K), $k\in\{10,20\}$. Due to space constraints, we included results with MAP and P@K only in the online appendix. Note that for test sets with sparse labels (MS MARCO development set and MS MARCO FarRelevant) we computed only MRR.
Figure 2. Zero-shot vs. fine-tuned performance on MS MARCO FarRelevant for a representative set of models.
Table 6. Model Ranking Performance on MS MARCO FarRelevant.
Retriever / Ranker zero-shot transferred fine-tuned
@@ 3.2. Setup
  • In each column we show a relative gain over models’ respective MaxP baseline. For LongP models, the gain is over MaxP (BERT). To reduce clutter we do not compute gains for FirstP models. Best numbers are in bold: Our results are averaged over three seeds. Unless specified explicitly, the backbone is BERT-base.
@@
3.3. Results
Our main experimental results for MS MARCO, TREC DL 2019-2021, and Robust04 are presented in Table 1. Table 6 and Fig. 2 shows results for MS MARCO FarRelevant.
We abbreviate names of several PARADE models: Note that PARADE Attn denotes a PARADE Attention model. The PARADE Transf prefix denotes PARADE Transformer models where an aggregator Transformer can be either trained from scratch (Transf-RAND-L2) or initialized with a pretrained model (Transf-PRETR-L6). L2 and L6 denote the number of aggregating layers (two and six, respectively).333 Note, however, that Transf-PRETR-L2 has only four attention heads. Unless explicitly specified, the backend Transformer model is BERT-base (Devlin et al., 2019).
To ease understanding and simplify presentation, we display key results for a representative sample of models in Fig. 1 (see § 1) and Fig. 2. Moreover, in Table 1 we present only a single aggregate number of all TREC DL query sets, which is obtained by combining all the queries and respective relevance judgements (i.e., we post an overall average rather than an average over the mean values for 2019, 2020, and 2020).
From Fig. 1, we learn that the maximum average gain over respective FirstP baselines is only 5% (when measured using MRR or NDCG@K). Gains are much smaller for a number of models, which even underperform their FirstP baselines on one or more dataset and some of these differences are statistically significant. In particular, this is true for CEDR-DRMM, CEDR-KNRM (MacAvaney et al., 2019), JINA (günther2023jina) and MOSAIC (Portes et al., 2023) on the MS MARCO development set.
@@ 3.3. Results
To understand if using BERT-base as backbone model negatively affected models’ ability to process and understand long contexts, we carried out a focused comparison of several backbone models, including ELECTRA (Clark et al., 2020) and DEBERTA (He et al., 2021). Here we use a PARADE Attention model, which achieved the largest average gain over FirstP for the BERT-base backend, as well as a MaxP model, which was extensively benchmarked in the past (Li et al., 2024; Dai and Callan, 2019; Zhang et al., 2021). Although prior work found ELECTRA to be a better backbone model in terms of absolute accuracy (Li et al., 2024; Zhang et al., 2021), we found no systematic evaluation of the relationship between a backbone model and achievable FirstP gains. Results in Tables 5 and 1 confirm overall superiority of both ELECTRA and DEBERTA over BERT-base. In that, DEBERTA models are nearly always more effective compared to ELECTRA models with biggest differences on Robust04. However, their relative effectiveness with respect to their FirstP baseline does not exceed that of BERT-base. The same is true for LongP models. Except Longformer they performed equally or worse compared to FirstP on 8 test sets out of 18. Moreover, all LongP models achieved lower average gains over FirstP (see the last column in Table 1). We conclude that to measure capabilities of chunk-and-aggregate model to understand and incorporate long context, it appears to be beneficial to use BERT-base instead of ELECTRA or DEBERTA.
We also use Table 5 to compare with prior art. We generally reproduce prior art, in particular, experiments by Li et al (Li et al., 2024), who invented PARADE models. Our ELECTRA-based models achieve higher NDCG@10 on TREC DL 2019-2020 and PARADE Attention models come very close, but they are about 3-5% worse compared to their PARADE Transformer on Robust04. At the same time, our DEBERTA-based PARADE Attention model achieves similar NDCG@20 scores.
Note that one should not expect identical results due to differences in training regimes and candidate generators. In particular, in the case of Robust04, Li et al (Li et al., 2024) use RM3 (BM25 with a pseudo-relevance feedback (Jaleel et al., 2004)), which is more effective than BM25 (Robertson, 2004) (which we use on Robust04).
Another important comparison point is Robust04 results by Zhang et al. (Zhang et al., 2021) who were able to reproduce original MaxP results by Dai and Callan (Dai and Callan, 2019), which used BERT-base as a backbone. In addition, they experimented with ELECTRA models “pre-finetuned” on MS MARCO. When comparing BERT-base results, Zhang et al. (Zhang et al., 2021) have the maximum relative gain of 6.6% compared to ours 3.3%. However, in absolute terms we got higher NDCG@20 for both FirstP and MaxP. Their MaxP (ELECTRA) has comparable performance to ours on TREC DL 2019-2020, but it is 2-4% better on Robust04. In turn, our MaxP (DEBERTA) is better by 2-6%. Although we do not always match prior art using the same backbone models, we generally match or outperform prior results, which, we believe, boosts the trustworthiness of our experiments.
Our analysis of position of relevance passages in MS MARCO as well as results by Hofstätter (Hofstätter et al., 2020) provide some evidence that limited benefits of long-context models are not due inability to process long context, but rather due to a positional bias of relevant passages, which tended to be among the first 512 document tokens (see Table 4). To further support this hypothesis, we carried out two sets of experiments using the MS MARCO FarRelevant collection, where a relevant passage is never in the first chunk. We carried out both the zero-shot experiment (evaluation of the model trained on MS MARCO) as well fine-tuning experiment using 50K MS MARCO FarRelevant queries. Because First perform poorly in this setting we use different baselines, namely, Longformer and MaxP models. For models with ELECTRA and DEBERTA backends we compare with MaxP (ELECTRA) and MaxP (DEBERTA), respectively. Otherwise, the baseline is MaxP (BERT). From Fig. 2 and Table 6, we make the following key observations:
@@
  • Not only positional bias diminished benefits of longer document contexts, but also lead to models’ overfitting to this bias and performing poorly in a zero-shot setting when the distribution of relevant passages changed substantially;
  • Although PARADE Transformer modelds were more effective than other models on standard collections, their advantage was small (a few %). In contrast, on MS MARCO FarRelevant, PARADE Transformer (ELECTRA) outperformed the next competitor Longformer by 8% and PARADE Max (ELECTRA)—an early chunk-and-aggregate approach—by as much as 23.8% (RQ2).

### v3

@@ 2.1 Related Work
2.2 Data
Our primary datasets include two MS MARCO Documents collections (v1 and v2) Bajaj et al. (2016); Craswell et al. (2020, 2022), Robust04 Voorhees (2004), and associated query sets. In addition, we created a collection MS MARCO FarRelevant by using passages and relevance judgments from the MS MARCO Passages collection.
Robust04 is a small collection of 0.5M documents that has a mixture of news articles and government documents some of which are quite long. Yet it has only a small number of queries (250), which makes it a challenging benchmark for training models in a low-data regime. Each query has a title and a description, which represent a brief information need and a more elaborate request (often a proper English prose), respectively. We use Robust04 in a cross-validation settings with folds established by Huston and Croft Huston and Croft (2014) provided via IR-datasets MacAvaney et al. (2021). All datasets are in English. Document and query statistics are summarized in Tables 2 and 3.
MS MARCO v1 was created from the MS MARCO reading comprehension dataset Bajaj et al. (2016) and it has two related collections: passages and documents. MS MARCO v1 comes with large query sets, which is particularly useful for training and testing models in the big-data regime. These query sets consist of question-like queries sampled from the Bing search engine log with subsequent filtering Craswell et al. (2020). Note that queries are not necessarily proper English questions, e.g., “lyme disease symptoms mood”, but they are answerable by a short passage retrieved from a set of about 3.6M Web documents Bajaj et al. (2016).
MS MARCO v1 test sets were created in two stages, where initially relevance judgements were created for the passage variant of the dataset. Then, document-level relevance labels were created by transferring passage-level relevance to original documents from which passages were extracted. To assess positional bias, we mapped relevant passages (from the MS MARCO Passage collection) to their positions in documents. Because document and passage texts were collected at different times this lead to some content divergence Craswell et al. (2020) and made exact mapping impossible: In particular, Hofstätter et al. 2020b were able to match only 32% of the passages:
We deemed such mapping insufficient: To obtain a more comprehensive mapping we resorted to approximate matching and were able to match about 85% of the passages. We manually inspected a sample of matched passages to ensure that the matching procedure was reliable. Moreover, the distribution of positions of relevant passages matched that of a related FIRA dataset Hofstätter et al. (2020b), where such information was collected by crowdsourcing. Positional bias information is summarized in Table 1.
Relevance labels in the training and development sets are “sparse”: There is about one positive example per query without explicit negatives. In addition to sparse relevance judgements—separated into training and developments subsets—there is a small number (98) of queries that have “dense” judgements provided by NIST assessors for TREC 2019 and 2020 deep learning (DL) tracks Craswell et al. (2020).
@@ 2.2 Data
The MS MARCO FarRelevant collection was created from the MS MARCO passage collection in such a way that each document contained exactly one relevant passage and this passage did not start before token 512 (see algorithm in the Appendix § B.1). Moreover, we created just a single relevant document for each training or testing query. MS MARCO FarRelevant is a variant of a the needle-in-the-haystack test Saad-Falcon et al. (2024); Zhu et al. (2024). It is designed to be textually similar to MS MARCO Documents but with different positional biases for relevant passages. Due MS MARCO having a non-commercial license, MS MARCO FarRelevant has the same licensing restriction.
Although we generated about 7K test queries and about 500K training queries, we used only 50K and 1K queries for fine-tuning and testing, respectively. On one hand, this was sufficient for accurate training and testing and, on the other hand, it reduced experimentation time and cost.
2.3 Overview of Selected Methods
Due to space constraints, a detailed description is given in the Appendix § A. In summary, all methods can be divided into split-and-aggregate (SplitP) methods and LongP methods that “natively” support longer documents inputs. SplitP use either simple aggregating operations (averaging, summing, taking the maximum) or an aggregator neural network. CEDR MacAvaney et al. (2019), PARADE Attention Li et al. (2024), and Neural Model 1 Boytsov and Kolter (2021) aggregate using simple neural networks, whereas PARADE Transformer models aggregator is a smaller Transformer Li et al. (2024).
We focused on cross-encoding rankers, which process queries concatenated with documents Nogueira and Cho (2019). As a reference point we also tested a bi-encoding E5-4K model, which had strong performance on LongEmbed benchmark with context sizes under 4K tokens Zhu et al. (2024). E5-4K was tested as a ranking model and only in the zero-shot mode (without fine-tuning).
Nearly all rankers use only BERT models (i.e., bi-directional encoder-only Transformers) and have in total 100M-200M parameters (see Table 6). In addition, inspired by a recent success of LLM-rankers Pradeep et al. (2023); Ma et al. (2023), we tested a much larger cross-encoding decoder-only (“causal”) Transformer model. Specifically we chose a 1B-parameter TinyLLAMA model due to its impressive performance for its relatively small size Zhang et al. (2024).
Table 4: Ranking Performance on MS MARCO, TREC DL, and Robust04.
Retriever / Ranker MS MARCO TREC DL Robust04 Avg. gain
dev (2019-2021) title description over FirstP
@@ 2.3 Overview of Selected Methods
retriever 0.312 0.629 0.428 0.402
FirstP (BERT) 0.394 0.632 0.475 0.527
FirstP (Longformer) 0.404 0.643 0.483 0.540
FirstP (ELECTRA) 0.417 0.662 0.492 0.552
FirstP (DEBERTA) 0.415 0.672 0.534 0.596
FirstP (Big-Bird) 0.408 0.656 0.507 0.560
FirstP (JINA) 0.422 0.654 0.488 0.532
FirstP (MOSAIC) 0.423 0.643 0.453 0.538
FirstP (TinyLLAMA) 0.395 0.615 0.431 0.473
FirstP (E5-4K) zero-shot 0.380 0.641 0.438 0.429
PARADE Attn 0.416a $(+5.5\%)$ 0.652a $(+3.1\%)$ 0.503a $(+5.7\%)$ 0.556a $(+5.6\%)$ +5.0%
LongP (TinyLLAMA) 0.402a $(+1.7\%)$ 0.608 $(-1.1\%)$ 0.452a $(+4.8\%)$ 0.505a $(+6.7\%)$ +$3.0\%$
LongP (E5-4K) zero-shot 0.353a $(-7.1\%)$ 0.649 $(+1.3\%)$ 0.439 $(+0.1\%)$ 0.434 $(+1.1\%)$$1.1\%$
    Statistical significant differences with respect to this baseline are denoted using the superscript superscript a.
    $p$-value threshold is 0.01 for an MS MARCO development collection and 0.05 otherwise.
3 Experiments
Table 5: Model Ranking Performance on MS MARCO FarRelevant.
Retriever / Ranker zero-shot fine-tuned
transferred
@@ 3 Experiments
[…]
LongP (E5-4K) 0.057ab $(-80.7\%)$ N/A (zero-shot only)
  • In each column we show a relative gain over models’ respective MaxP baseline. For LongP models, the gain is over MaxP (BERT). Best numbers are in bold: Our results are averaged over three seeds. Unless specified explicitly, the backbone is BERT-base.
@@
3.1 Setup
We trained each cross-encoding ranking model using three seeds, except the bi-encoder model E5 Zhu et al. (2024), which was evaluated only in the zero-shot mode. To compute statistical significance, we averaged query-specific metric values over these seeds. Due to space constraints, additional experimental details are provided in the Appendix § B.2. Moreover, in the main part of the paper we only show results for the mean reciprocal rank (MRR) and the non-discounted cumulative gain at rank $k$ (NDCG@K). Additional precision-related metrics are computed in the Appendix (see § B.5).
3.2 Results
Our main experimental results for MS MARCO, TREC DL 2019-2021, and Robust04 are presented in Table 4. Table 5 and Fig. 2 show results for MS MARCO FarRelevant. In the Appendix (see B.4) we show that we can match or outperform key prior results, which, we believe, boosts the trustworthiness of our experiments.
We abbreviate names of several PARADE models: Note that PARADE Attn denotes a PARADE Attention model. The PARADE Transf or P. Transf prefix denotes PARADE Transformer models where an aggregator Transformer can be either trained from scratch (Transf-RAND-L2) or initialized with a pretrained model (Transf-PRETR-L6). L2 and L6 denote the number of aggregating layers (two and six, respectively).333 Note, however, that Transf-PRETR-L2 has only four attention heads.
Unless explicitly specified, the backbone Transformer model for SplitP methods is BERT-base Devlin et al. (2019). Although using other backbones such as ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021) can improve an overall accuracy, we observe a bigger gain compared to a FirstP baseline when we use BERT-base (see § B.4 in the Appendix).
To ease understanding and simplify presentation, we display key results for a representative sample of models in Fig. 1 and Fig. 2 (in § 1). Moreover, in Table 4 we present only a single aggregate number for all TREC DL query sets, which is obtained by combining all the queries and respective relevance judgements (i.e., we post an overall average rather than an average over the mean values for 2019, 2020, and 2020).
From Fig. 1 and Table 4 we learn that the maximum average gain over respective FirstP baselines is only 5% (when measured using MRR or NDCG@K). Gains are much smaller for a number of models, which even underperform their FirstP baselines on one or more dataset and some of these differences are statistically significant. In particular, this is true for CEDR-DRMM, CEDR-KNRM MacAvaney et al. (2019), JINA Günther et al. (2023) and MOSAIC Portes et al. (2023) on the MS MARCO development set.
@@ 3.2 Results
Our analysis of position of relevance passages in MS MARCO as well as results by Hofstätter et al. 2020b provide strong evidence that limited benefits of long-context models are not due inability to process long context, but rather due to a positional bias of relevant passages, which tended to be among the first 512 document tokens (see Table 1).
To further support this hypothesis, we carried out two sets of experiments using the MS MARCO FarRelevant collection, where a relevant passage was never in the first chunk. We carried out both the zero-shot experiment (evaluation of the model trained on MS MARCO) as well fine-tuning experiment using 50K MS MARCO FarRelevant queries. Because FirstP models perform poorly in this setting we use different baselines, namely, Longformer and MaxP models. For models with ELECTRA and DEBERTA backbones we compare with MaxP (ELECTRA) and MaxP (DEBERTA), respectively. Otherwise, the baseline is MaxP (BERT). From Fig. 2 and Table 5, we make the following key observations:
  • The FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning modes (RQ3). Surprisingly, E5-4K performance is also at a random-baseline level despite its competitive performance on LongEmbed benchmark Zhu et al. (2024), MS MARCO, and Robust04 (see Table 4);
@@
  • Not only positional bias diminished benefits of supporting longer document contexts, but it also lead to model overfitting to the bias and performing poorly in a zero-shot setting when the distribution of relevant passages changed substantially;